Q.solve the following:


Sin-1 + cos -11-4x2 = /3


(square root is on whole 1-4x2)

Grade:11

1 Answers

jitender lakhanpal
12 years ago

Dear SHOBHIT,

AS WE KNOW

  COS-1 X =SIN- 1  √  1-X2

SO WE CAN WRITE  COS-1√ 1-4X2  =SIN-1( 2X)

                                                  =SIN-1-( 2X)

WE GET SIN-1( 2X)+SIN-1( X)=    ∏/3----EQUATION 1  AND      SIN-1( 2X)-SIN-1( X)=    ∏/3   ----EQUATION 2

SOLVING EQUATION 1 

SIN-1( 2X) = ∏/3- SIN-1( X)

2X = SIN∏/3 COS(SIN-1( X) ) - COS ∏/3 SIN(SIN-1( X))

2X = √3/2  √  1-X2 -(1/2)X

FROM THE ABOVE EQUTION WE GET X = -1/2

AND SIMILARLY WE GET OTHER VALUES BY SOLVING THE EQUTION 2 



We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar: Click here to download the toolbar.. 

JITENDER

ASKIITIANS EXPERT